P years ago, a man’s age was (q+r)times that of his son who was then r years old.How old is the man now?
Answer by mananth(16946)

P years ago, a man’s age was (q+r)times that of his son who was then r years old.How old is the man now?
Let man's age now be x
p years ago
he was (x-p) years
(x-p) = (q+r)r
x= r^2 +qr +p